OpenStudy (anonymous):

What is the solution to the equation 3/4(4c + 16) = 2c + 9 ? c =

OpenStudy (anonymous):

\[3(4c+16) = 4(2c+9)\]

OpenStudy (anonymous):

are you able to work that out now?

OpenStudy (emmamink):

12c + 12 = 2c +9 3=-10c -3/10

OpenStudy (emmamink):

-3/10 is the answer

OpenStudy (anonymous):

3 times 4c + 16 does not equal 12c + 12, try again

OpenStudy (emmamink):

um 3/4c does

OpenStudy (emmamink):

ohh opps i screwed up

OpenStudy (emmamink):

actually no i didnt

OpenStudy (emmamink):

-3/10 stands

OpenStudy (anonymous):

EmmaMink watch this:\[3/4 (4c+16)= 3(4c+16)/4\]

OpenStudy (emmamink):

where did u get that from?

OpenStudy (anonymous):

@BPDlkeme234 that is not the math problem

OpenStudy (anonymous):

that means you can simply cross multiply to get 4* (2c + 9)

OpenStudy (emmamink):

i know BPDIkeme234 is crazy the answer is -3/10 don't listen to him

OpenStudy (emmamink):

So, help?!?!?!, that is the answer

OpenStudy (anonymous):

No that is not the answer

OpenStudy (emmamink):

-3/10? yes it is

OpenStudy (anonymous):

right so, you end up with 3(4c+16) = 4(2c+9)

OpenStudy (anonymous):

multiplying this out you get what?

OpenStudy (anonymous):

3c+12=8c+36

OpenStudy (emmamink):

oops -3 that is the answer. And symbolab.com is a great equation solver! so -3 is the answer

OpenStudy (anonymous):

I got 12c+48=8c + 36

OpenStudy (emmamink):

-3 people!!!!

OpenStudy (anonymous):

now, 12c - 8c = 36 -48

OpenStudy (anonymous):

i.e 4c = -12

OpenStudy (anonymous):

c = -12/4

OpenStudy (emmamink):

the answer is -3!!

OpenStudy (anonymous):

how did you get 48? 3/4*16=12

OpenStudy (anonymous):

you guessed it @EmmaMink

OpenStudy (emmamink):

ok so we agree? -3 is the answer.

OpenStudy (anonymous):

yes @EmmaMink the answer is -3!

OpenStudy (emmamink):

ok

OpenStudy (emmamink):

btw u a boy or girl

OpenStudy (anonymous):

I would like to reply to @help?!?!?! who seems to be confused

OpenStudy (anonymous):

I am a boy (man actually)

OpenStudy (emmamink):

oh ok

OpenStudy (anonymous):

I also work as a freelance maths tutor

OpenStudy (emmamink):

kewl!

OpenStudy (anonymous):

so ask me any questions, I love the challenge!

OpenStudy (anonymous):

Thank you